What is the volume of a solution that contains 0.50 mole of 3.6 M KMnO4?
Answer #1

Let volume is V liters.

Molarity = Moles/Volume of solution in liters

3.6 = 0.50/V

V = 0.50/3.6 = 0.1389 liters

OR

V = 138.9 ml ......Answer